Circular No. 7999

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2002hb, 2002hc, 2002hd Three apparent supernovae have been discovered. SN 2002hb was

found by W. M. Wood-Vasey, G. Aldering, and P. Nugent on unfiltered NEAT images taken with the Palomar 1.2-m Schmidt telescope (cf. IAUC 7996). SNe 2002hc and 2002hd were found on unfiltered KAIT CCD images taken by LOTOSS (cf. IAUC 7906) and reported by B. Beutler, P. Martin, M. Papenkova, and A. V. Filippenko. Additional magnitudes from images taken by the respective groups for each object: SN 2002hb, Sept. 3, [20.5 (S/N = 5); Oct. 3, 18.9; 21, 19.3. SN 2002hc in NGC 2559, Apr. 9, [19.0:; Oct. 24, 17:. SN 2002hd in MCG –01-23-8, Apr. 9, [19.0:; Oct. 25, 18:.

SN  2002 UT  α_{2000}  δ_{2000}  Mag.  Offset
2002hb Sept. 11.35 22^{h}56^{m}03.^{s}58 +13^{o}37′31′′.0 18.7 0′′.7 S
2002hc Oct. 23  8 17 06.05 −27 27 27.0 17  1′′.1 E, 7′′.2 S
2002hd Oct. 24  8 54 04.68 − 7 10 59.5 18  19′′.5 W, 21′′.5 S

M. Hamuy, Carnegie Observatories, reports that a spectrum (range 360–900 nm) of SN 2002hc, obtained on Oct. 25.34 UT with the Las Campanas Baade 6.5-m telescope (+ LDSS2), shows it to be a type-II supernova. A broad Hα line is seen in emission with no significant absorption counterpart. The spectrum is very similar to that of the ‘linear supernova’ 1979C when 43 days past maximum (Filippenko 1997, ARAA 35, 309).

SUPERNOVAE 2002gy AND 2002ha Hamuy also reports that a spectrum of SN 2002ha (cf. IAUC 7997),

obtained as above on Oct. 25.03 UT, shows it to be a type-Ia supernova ∼ 3 days before maximum light, very similar to SN 1994D. A narrow interstellar Na I D absorption line at the redshift of the host galaxy can be seen with an equivalent width of 0.1 nm. Hamuy adds that a spectrum of SN 2002gy (cf. IAUC 7996), obtained on Oct. 25.21, bears resemblance to the type-Ib/c SN 1999ex. A prominent absorption at rest wavelength 614 nm can be seen, which could be due to Si II 635.5-nm or high-velocity Hα. A narrow interstellar Na I D absorption line at the redshift of the host galaxy is present with an equivalent width of 0.32 nm, which suggests significant reddening.

2002 October 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green